# Angular momentum of vacuum bubbles in a first-order phase transition

### v1

@@ 4.2.5 Numerical scheme
$\tilde{\eta}_{i}=0.1-10^{-3},\tilde{\eta}_{*}=0.1,c_{s\rm,D*}^{2}=2\times 10^{-4}$ $\tilde{\eta}_{i}=0.1-10^{-4},\tilde{\eta}_{*}=0.1,c_{s\rm,D*}^{2}=2\times 10^{-4}$
Refer to caption Refer to caption
$\tilde{\eta}_{i}=1-10^{-3},\tilde{\eta}_{*}=1,c_{s\rm,D*}^{2}=2\times 10^{-4}$ $\tilde{\eta}_{i}=1-10^{-4},\tilde{\eta}_{*}=1,c_{s\rm,D*}^{2}=2\times 10^{-4}$
Figure 11: The density and velocity transfer functions, approximating $c_{s\rm,D}^{2}$ via Eq. (4.88), when the width of $c_{s\rm,D}^{2}$ (i.e. changing $\tilde{\eta}_{i}$) or the sound speed squared at percolation (i.e. changing $c_{s\rm,D*}^{2}$) is changed. In the last row, we also changed the duration of the FOPT, by a factor of 10, while still keeping the width $\tilde{\eta}_{*}-\tilde{\eta}_{i}$.
Modes with sufficiently long wavelengths are marginally affected by the sudden drop in $c_{s\rm,D}^{2}$, and thus the amplitude of the density (velocity) transfer function for such modes approaches $4$ ($4/\sqrt{3}$). The transition to the asymptotic values of the transfer function for $k/\mathcal{H}_{\rm c}\gg 1$ must be continuous, as shown by the transfer functions in Fig. 11. We generated those transfer functions, under the assumption that we approximate the sound speed squared as a piecewise linear function in rescaled conformal time. The range of $k$-modes that connect the amplitudes of the transfer functions for small $k$ and large $k$, can be obtained as follows. From physical grounds, modes that satisfy
$\frac{k}{\mathcal{H}_{\rm c}}\gtrsim\frac{1}{\tilde{\eta}_{*}-\tilde{\eta}_{i}}$ (4.98)
will have sufficient temporal resolution to be affected by the decrease in $c_{s\rm,D}^{2}$ around the percolation time. The corresponding density modes will undergo enhancement to $4/(3c_{s\rm,D*}^{2})^{1/4}$, up to a certain $k$-mode where the next-to-leading order term in Eq. (4.93) is subdominant; an analogous statement can also be said for the velocity modes. Meanwhile, an upper limit in $k/\mathcal{H}_{\rm c}$, beyond which the density and velocity transfer functions are guaranteed to be quenched to their asymptotic values, is
$\frac{k}{\mathcal{H}_{\rm c}}<\frac{1}{(3c_{s\rm,D*}^{2})^{3/2}\left(\tilde{\eta}_{*}-\tilde{\eta}_{i}\right)}\,.$ (4.99)
In Fig. 11, vertical dashed lines mark the range of modes identified by Eqs. (4.98) and (4.99).
The approximate treatment of assuming a linear drop in $c_{s\rm,D}^{2}$ allows us to identify the factors that affect the transfer functions, namely: the period $\tilde{\eta}_{*}-\tilde{\eta}_{i}$ when the drop in $c_{s\rm,D}^{2}$ occurs; and the sound speed squared $c_{s\rm,D*}^{2}$ at the percolation time. We demonstrate the effect of these quantites on the transfer functions in Fig. 11. In the top row we changed $\tilde{\eta}_{*}-\tilde{\eta}_{i}$, effectively changing the width of the dip in $c_{s\rm,D}^{2}$ (across columns). As expected, reducing the width simply moves the locations of the vertical lines that mark those modes that suffer enhancement in the density. For the middle row, we changed $c_{s\rm,D*}^{2}$ (from top to middle row). Decreasing $c_{s\rm,D*}^{2}$ increases/decreases the quenched values of the density/velocity transfer functions in the large $k/\mathcal{H}_{\rm c}$ limit. The bottom row features transfer functions for a longer FOPT duration. In keeping the width and $c_{s\rm,D*}^{2}$ the same as in the middle row, we do not see any noticeable change in behavior in the transfer functions. Finally, in Fig. 10, the horizontal dashed lines demonstrate that our estimates for the amount of enhancement/suppression in the density/velocity transfer functions match the results of numerical calculation.
@@ 5.1 Numerical values of physical quantities
The coefficient $\mathcal{C}_{*}$, expressed as an integral in Eq. (3.43) and evaluated at the percolation time, contains explicit dependence on the primordial power spectrum of curvature perturbations, and the transfer functions for all $k$ modes. However, from a practical standpoint, only a finite number of modes can be calculated numerically. We have chosen to solve the perturbation equations up to the percolation time, for $N=1000$ modes between $10^{-3}<k/\mathcal{H}_{\rm c}<5/(H_{*}R_{*})$ that are uniformly spaced in logarithm, and construct a table of values for $\delta_{\rm D,*}$, $\hat{\psi}_{\rm D,*}$, $\Phi_{*}$, and $-k^{2}\gamma_{*}$ versus $k/\mathcal{H}_{*}$. Then, for an arbitrary $k/\mathcal{H}_{*}$, we can evaluate these perturbations from the table by interpolation. We then perform the $\mathcal{C}_{*}$ integration using a general purpose Monte Carlo integrator, and by setting the cutoff at
@@
which corresponds to a mode whose wavelength matches the FV bubble diameter. We do not consider modes that are shorter than $k_{\rm cut}$, since they can resolve between false and true vacuum regions, such that the dark plasma can no longer be considered as a single entity.
@@
Figure 12: $\mathcal{C}_{*}$ and $v_{\rm eff,*}$ versus $\tilde{x}_{0,*}$ for the FOPT benchmark cases BP-2 (panel a) and BP-7 (panel b) in Fig. 3. The vertical line marks the size of the FV bubble $R_{\rm FV,*}$ relative to Hubble constant at percolation.
In Fig. 12 we show the dependence of $\mathcal{C}_{*}$ and $v_{\rm eff,*}$ on the bubble size relative to Hubble in red and blue, respectively. The two panels feature BP-2 and BP-7. The dotted lines, corresponding to the SM approximation, use the exact transfer functions in Eqs. (4.14), to calculate $\mathcal{C}_{*}$ and $v_{\rm eff,*}$, with integration up to $k=\pi/R_{*}$. The vertical line marks the size of the FV bubble $R_{\rm FV,*}$ relative to Hubble, associated with the prescription in Eq. (4.83). As we can see in Fig. 12, the difference between the dark sector $v_{\rm eff,*}$ and the SM approximation is at most an order 1 factor. The drop in both $\mathcal{C}_{*}$ and $v_{\rm eff,*}$ for $\tilde{x}_{0,*}<\sqrt{3}H_{*}R_{\rm FV,*}$ is associated with the fact that $k_{\rm cut}$ is above the largest $k$ mode available in the interpolation table; whenever we encounter this case, we extrapolate the transfer functions to zero. For the benchmark points in Fig. 12, the relative size of the FV bubble, $\mathcal{C}_{*}$, $v_{\rm eff,*}$, and $s_{\rm rms,*}$, can be found in Table 1. We find that the FV bubbles are much smaller than the Hubble horizon and the effective equatorial velocity is much smaller than the speed of light.
@@ 5.2 Estimates and scans over the physical region
$\rho_{\rm tot,*}=\frac{\pi^{2}}{30}g_{\rho,{\rm SM}}(T_{*})\frac{T_{*}^{4}}{r_{T\rm,c}^{4}}+\frac{\pi^{2}}{30}g_{\rho}T_{*}^{4}-(1-F_{*})L_{\rm c}=\frac{3H_{*}^{2}}{8\pi G_{\rm N}}\,.$ (5.6)
@@
$\boxed{s_{\rm rms,*}\approx\left(\frac{2}{3}\right)^{1/2}\left(\frac{\beta_{*}/H_{*}}{v_{\rm w,*}}\right)^{2}~{}\frac{g_{\rho,{\rm SM}}(T_{*})}{g_{\rm*}}r_{T\rm,c}^{-4}\left[1-(1-F_{*})\frac{30}{\pi^{2}g_{\rho}}\frac{L_{\rm c}}{T_{\rm c}^{4}}\left(\frac{T_{\rm c}}{T_{*}}\right)^{4}\right]^{-1}v_{\rm eff,*}}\,.$ (5.7)
@@
Figure 13: Contour plots of $s_{\rm rms,*}/(v_{\rm eff,*}/A_{\rm s})$ for $A=10^{-1},10^{-2},10^{-3},10^{-4}$, for fixed $\lambda/\lambda_{0}=1.42$, and $T_{\rm c}=10\,\mathrm{MeV}$, $r_{T\rm,c}=0.4$.
Note that the explicit dependence of $s_{\rm rms,*}$ on $v_{\rm w}$ means that the spin depends on the exponent $k$ in in Eq. (4.23). For BP-2 and BP-7, we have checked that for $k$ values from 2 to 8, the spin gradually decreases by a factor roughly equal to $\left[1-(T_{*}/T_{\rm c})^{8}\right]^{2}/\left[1-(T_{*}/T_{\rm c})^{2}\right]^{2}$. Note that $\beta_{*}/H_{*}$, $v_{\rm eff,*}$, and $T_{*}$ are marginally sensitive to $k$. One can show that the term in square brackets is significant when $L_{\rm c}/T_{\rm c}^{4}$ hits the maximum possible value allowed by the physical conditions; otherwise, for $L_{\rm c}/T_{\rm c}^{4}\ll g_{\rho}\pi^{2}/30$, the term in square brackets is simply $F_{*}$. On the other hand, we typically find that $v_{\rm eff,*}/A_{\rm s}$ is in the range $\sim 10^{-3}$ to $\sim 10^{-2}$. We have found that the value of $v_{\rm eff,*}/A_{\rm s}$ is sensitive to the spectral index $n_{\rm s}$ of the primordial curvature power spectrum. In the case where $n_{\rm s}=1$, $v_{\rm eff,*}/A_{\rm s}$, and thus the spin value, is enhanced by roughly an order of magnitude.
@@
Table 4: Spin and other physical properties of typical astrophysical objects
We can see from Fig. 13 that a wide range of $s_{\rm rms,*}/(v_{\rm eff,*}/A_{\rm s})$ values can be obtained, even for fixed $A$ and $\lambda$, and fixed $T_{\rm c}$ and $r_{T\rm,c}$. Equation (5.7) explains this behavior because the percolation temperature and $\beta_{*}/H_{*}$ vary widely within the physical region. Furthermore, we see the effect of changing $T_{\rm c}$ and $r_{T\rm,c}$ in the contours of $s_{\rm rms,*}/(v_{\rm eff,*}/A_{\rm s})$ in Fig. 14. Panel (d) shows the scan for the default choice $T_{\rm c}=10\,\mathrm{MeV},r_{T\rm,c}=0.4$. Panels (a) and (b) show the effect of changing $T_{\rm c}$, which increases in value for larger $T_{\rm c}$ due to the increase in $g_{\rho,{\rm SM}}$, consistent with Eq. (5.7). On the other hand, comparing panels (c) and (d) illustrates the effect of changing $r_{T\rm,c}$, where a smaller $r_{T\rm,c}$ increases the overall $s_{\rm rms,*}/(v_{\rm eff,*}/A_{\rm s})$.
@@
Figure 14: Contour plots of $s_{\rm rms,*}/(v_{\rm eff,*}/A_{\rm s})$ for $A=10^{-3}$, changing $T_{\rm c}$ (top) and $r_{T\rm,c}$ (bottom).
@@ 5.3 Benchmarks
To obtain a general idea of the prediction for the RMS spin value of FV bubbles, we consider a collection of FOPT scenarios that will serve as our benchmark points. We require that these FOPT scenarios are, at least, physically viable. Then we calculate the transfer functions for these benchmarks, in order to eventually obtain $v_{\rm eff,*}$. One criterion we impose on these benchmarks is $T_{*}/T_{\rm c}\leq 0.95$, which guarantees that the effective potential will have a sufficiently high barrier between the false vacuum and the true vacuum.
@@
As usual, we had taken $A=10^{-1},10^{-2},10^{-3},10^{-4}$; then for each $A$, we have selected $\lambda/\lambda_{0}(A)=1.42,2.34,2.94,5$. For each pair of $(A,\lambda)$, we scan over the physical region in the $L_{\rm c}/T_{\rm c}^{4}-x_{C}$ plane and obtain physical quantities such as $\alpha_{*}$, $T_{*}$, and $s_{\rm rms,*}/(v_{\rm eff,*}/A_{\rm s})$. We choose a random collection of FOPT scenarios and solve the perturbation equations to construct a table of transfer functions versus $k/\mathcal{H}_{\rm c}$. This allows us to calculate $\mathcal{C}_{*}$ and the effective equatorial velocity $v_{\rm eff,*}/A_{\rm s}$, and finally $s_{\rm rms,*}$. The final benchmark points are presented in Table 1, with the rows organized in three groups. The first group (from the top) has parameters that specify the effective potential, critical temperature, and initial temperature ratio. The second group has quantities that can be obtained solely by tracking the evolution of the background cosmology during the phase transition. The third group has quantities that are relevant in inferring the features of the transfer functions, as well as quantities that can be obtained after solving the perturbation equations.
The benchmarks selected have notable features. Most of the benchmark points have $r_{T\rm,c}=0.4$. BP-1 and BP-3 have $T_{\rm c}=100\,\mathrm{GeV}$, but correspond to different values of $A$. BP-1 and BP-2 have the same set of FOPT parameters $\{A,x_{C},\lambda,L_{\rm c}/T_{\rm c}^{4}\}$ but BP-2 has a lower $T_{\rm c}=10\,\mathrm{MeV}$. BP-5 corresponds to an FOPT with $T_{\rm c}=1\,\mathrm{MeV}$. BP-8 has $T_{\rm c}=10\,\mathrm{keV}$ and $r_{T,\rm c}=0.1$; it also gives the largest spin, while BP-6 features an FOPT with the smallest RMS spin among the benchmarks. Among the benchmarks, BP-7 has the lowest $T_{*}/T_{\rm c}$, so that its rescaled elapsed time since the critical point $\tilde{\eta}_{*}$ is the longest among the benchmarks.
Comparing BP-1 and BP-2, which only differ in $T_{\rm c}$, the change in RMS spin can be attributed to both the scaling in $(\beta/H_{*})/v_{\rm w,*}$, and in $g_{\rho,{\rm SM}}$. In particular, while $g_{\rho,{\rm SM}}$ decreased by around a factor of 10 from BP-1 to BP-2, $(\beta/H_{*})/v_{\rm w,*}$ increased by around a factor of 1.7; overall, the quantity $s_{\rm rms,*}/(v_{\rm eff,*}/A_{\rm s})$ only decreased by a factor of around 3. We emphasize that $(\beta/H_{*})/v_{\rm w,*}$ is roughly the inverse of the bubble radius relative to the Hubble scale, according to Eq. (5.4), i.e. a larger $\sqrt{3}H_{*}R_{*}$ means a smaller $(\beta_{*}/H_{*})/v_{\rm w,*}$. A similar trend can be seen for BP-3 and BP-5, both with $A=10^{-2}$, where a decrease of $\mathcal{O}(10)$ in $g_{\rho,{\rm SM}}$ and an increase of $\mathcal{O}(1)$ in $(\beta/H_{*})/v_{\rm w,*}$ only leads to a decrease of $\mathcal{O}(1)$ in $s_{\rm rms,*}/(v_{\rm eff,*}/A_{\rm s})$. BP-4 and BP-5 have $T_{\rm c}$ at the MeV scale. Their $g_{\rho,{\rm SM}}$ are both $\mathcal{O}(10)$, since the corresponding temperature in the SM sector is already below the QCD scale, where there is a marked decrease in the SM degrees of freedom. Analogously, their respective $(\beta_{*}/H_{*})/v_{\rm w,*}$ is $\mathcal{O}(10^{3})$. The decrease in $s_{\rm rms,*}/(v_{\rm eff,*}/A_{\rm s})$ from BP-4 to BP-5 is mainly attributed to the increase in $r_{T\rm,c}$ by a factor of 4, and is consistent with the scaling of $s_{\rm rms,*}/(v_{\rm eff,*}/A_{\rm s})$ as $r_{T\rm,c}^{-4}$.
BP-6 provides the smallest RMS spin value in the whole table, and this is mainly due to its $(\beta_{*}/H_{*})/v_{\rm w,*}$ being the lowest in the whole table; equivalently, BP-6 has the largest $\sqrt{3}H_{*}R_{*}$ value in the table. In contrast, BP-8 has the smallest $\sqrt{3}H_{*}R_{*}$, so it has the largest $(\beta_{*}/H_{*})/v_{\rm w,*}$ ratio. Furthermore, $r_{T\rm,c}=0.1$ for BP-8, making its spin even larger. In principle, the RMS spins can be made higher by lowering the temperature ratio $r_{T\rm,c}$; as a trade-off, the phase transition strength becomes weaker.

### v2

@@ 4.2.5 Numerical scheme
$\Delta\tilde{\eta}_{*}=10^{-3},\tilde{\eta}_{*}=0.1,c_{s\rm,D*}^{2}=-60$ $\Delta\tilde{\eta}_{*}=10^{-4},\tilde{\eta}_{*}=0.1,c_{s\rm,D*}^{2}=-60$
Figure 10: The density and velocity transfer functions, approximating $c_{s\rm,D}^{2}$ via Eq. (4.83), when the width of the dip in $c_{s\rm,D}^{2}$ (top) or the sound speed squared at percolation (bottom) is changed. The horizontal dashed lines correspond to the amplitude of the density and velocity transfer functions for perturbations in a pure radiation background. The vertical dashed lines are the critical Fourier scales where the exponential enhancement for the density and velocity perturbations becomes relevant. The dot dashed vertical line is the scale $1/\Delta\tilde{\eta}_{*}$, and the dotted line marks the Fourier scale beyond which the term of order $|c_{s\rm,D*}^{2}|/\Delta\chi_{*}$ becomes subdominant.
@@ 5.1 Numerical values of physical quantities
The coefficient $\mathcal{C}_{*}$, expressed as an integral in Eq. (3.43) and evaluated at the percolation time, contains explicit dependence on the primordial power spectrum of curvature perturbations, and the transfer functions for all $k$ modes. We perform the $\mathcal{C}_{*}$ integration using a general purpose Monte Carlo integrator, where we evaluate the integrand using a lookup table for the transfer functions covering the range $10^{-3}<k/\mathcal{H}_{\rm c}<k_{\rm max}/\mathcal{H}_{\rm c}$, and setting the cutoff scale in the integral at
@@
which corresponds to a mode whose wavelength matches the FV bubble diameter. We do not consider modes that are shorter than $k_{\rm cut}$, since they can resolve the false and true vacuum regions, and the dark plasma can no longer be considered as a single entity. Furthermore, the cutoff in $k$ can be justified by looking at the volume average, within a sphere of radius $R$, of the mean square density contrast in position space:
$\langle\delta^{2}(\vec{r})\rangle_{V}$ $\equiv$ $\frac{1}{V_{R}}\int_{R}d^{3}\vec{r}~\langle\delta^{2}(\vec{r})\rangle$ (5.4)
$=$ $\int\frac{d^{3}\vec{k}}{(2\pi)^{3}}\frac{d^{3}\vec{k}^{\prime}}{(2\pi)^{3}}~\langle\delta_{\vec{k}}\delta_{\vec{k}^{\prime}}\rangle~\frac{1}{V_{R}}\int_{R}d^{3}\vec{r}~\exp\left[-i\left(\vec{k}-\vec{k}^{\prime}\right)\cdot\vec{r}\right]$ (5.5)
$=$ $\lim_{k_{\rm th}\rightarrow\infty}\int_{0}^{k_{\rm th}}\frac{dk}{k}~\frac{k^{3}P_{\mathcal{R}}(k)}{2\pi^{2}}\left|U_{\delta}(k)\right|^{2}\,.$ (5.6)
Figure 11: Plot of the mean square density contrast as a function of the integral cutoff $k_{\rm th}/\mathcal{H}_{\rm c}$, for BP-2 (left) and BP-4 (right). The red dashed curves represent the approximate mean squared density contrast, which is obtained by replacing the transfer function in Eq. (5.6) with Eq. (4.93). The dashed and dot-dashed vertical lines respectively indicate $k_{\rm cut}/\mathcal{H}_{\rm c}$, and the value of $k/\mathcal{H}_{c}$ at which the RMS density contrast is 0.4.
As Fig. 11 shows, $\langle\delta^{2}\rangle_{V}$ blows up if $k_{\rm th}$ is taken to infinity. A finite cutoff will, of course, lead to a finite value of $\langle\delta^{2}\rangle_{V}$, and we are interested in knowing the cutoff, $k_{\rm th}/\mathcal{H}_{\rm c}$, where the mean square density contrast is at most the square of the threshold density contrast for PBH formation, which we take to be $\delta_{\rm th}^{2}\simeq 0.16$. For both BP-2 and BP-4, we find that $k_{\rm th}/\mathcal{H}_{\rm c}$ is smaller than $k_{\rm cut}/\mathcal{H}_{\rm c}$.
@@
Figure 12: $\mathcal{C}_{*}$ and $v_{\rm eff,*}$ versus $\tilde{x}_{0,*}$ for the FOPT benchmark cases BP-2 (panel a) and BP-4 (panel b) in Fig. 3. The vertical line marks the size of the FV bubble $R_{\rm FV,*}$ relative to Hubble constant at percolation.
In Fig. 12 we show the dependence of $\mathcal{C}_{*}$ and $v_{\rm eff,*}$ on the bubble size relative to Hubble in red and blue, respectively. The two panels feature BP-2 and BP-4. The dotted lines, corresponding to the SM approximation, use the exact transfer functions in Eqs. (4.14), to calculate $\mathcal{C}_{*}$ and $v_{\rm eff,*}$, with integration up to $k=\pi/R_{*}$. The vertical line marks the size of the FV bubble $R_{\rm FV,*}$ relative to Hubble, associated with the prescription in Eq. (4.90). As we can see in Fig. 12, the dark sector $v_{\rm eff,*}$ and the SM approximation can differ by an order of magnitude. The drop in both $\mathcal{C}_{*}$ and $v_{\rm eff,*}$ for $\tilde{x}_{0,*}<\sqrt{3}H_{*}R_{\rm FV,*}$ is associated with the fact that $k_{\rm cut}$ is above the largest $k$ mode available in the interpolation table; whenever we encounter this case, we extrapolate the transfer functions to zero. For the benchmark points in Fig. 12, the relative size of the FV bubble, $\mathcal{C}_{*}$, $v_{\rm eff,*}$ and $s_{\rm rms,*}$, can be found in Table 1. We find that the FV bubbles are much smaller than the Hubble horizon and the effective equatorial velocity is much smaller than the speed of light.
@@ 5.2 Estimates and scans over the physical region
$\rho_{\rm tot,*}\simeq\frac{\pi^{2}}{30}g_{\rho,{\rm SM}}(T_{*})\frac{T_{*}^{4}}{r_{T,*}^{4}}+\frac{\pi^{2}}{30}g_{\rho}T_{*}^{4}-(1-F_{*})\Delta\rho(T_{*})=\frac{3H_{*}^{2}}{8\pi G_{\rm N}}\,.$ (5.9)
@@
$\boxed{s_{\rm rms,*}\approx\left(\frac{2}{3}\right)^{1/2}\left(\frac{\beta_{*}/H_{*}}{v_{\rm w,*}}\right)^{2}\left\{1+\frac{g_{\rho,{\rm SM}}(T_{*})}{g_{\rho}}r_{T\rm,*}^{-4}\left[1-(1-F_{*})\frac{30}{\pi^{2}g_{\rho}}\frac{\Delta\rho(T_{*})}{T_{\rm c}^{4}}\frac{T_{\rm c}^{4}}{T_{*}^{4}}\right]^{-1}\right\}v_{\rm eff,*}\,.}$
Note that the explicit dependence of $s_{\rm rms,*}$ on $v_{\rm w}$ means that the spin depends on the exponent $n_{\rm w}$ in Eq. (4.23). For BP-2 and BP-4, we have checked that for $n_{\rm w}$ values from 2 to 8, the spin gradually decreases by a factor roughly equal to $\left[1-(T_{*}/T_{\rm c})^{8}\right]^{2}\left[1-(T_{*}/T_{\rm c})^{2}\right]^{-2}$. Note that $\beta_{*}/H_{*}$, $v_{\rm eff,*}$, and $T_{*}$ are marginally sensitive to $n_{\rm w}$. One can show that the term in square brackets is significant when $\Delta\rho(T_{*})/T_{\rm c}^{4}$ hits the maximum possible value allowed by the physical conditions; otherwise, for $\Delta\rho(T_{*})/T_{\rm c}^{4}\ll g_{\rho}\pi^{2}/30$, the term in square brackets is simply unity. On the other hand, we typically find that $v_{\rm eff,*}/A_{\rm s}$ is in the range $\sim 10^{-5}$ to $\sim 10^{-3}$. We have found that the value of $v_{\rm eff,*}/A_{\rm s}$ is sensitive to the spectral index $n_{\rm s}$ of the primordial curvature power spectrum. In the case where $n_{\rm s}=1$, $v_{\rm eff,*}/A_{\rm s}$, and thus the spin value, is enhanced by roughly an order of magnitude.
@@
Figure 13: Contours of constant $\log_{10}\left[s_{\rm rms,*}/(v_{\rm eff,*}/A_{\rm s})\right]$ in steps of 0.5 for $A=10^{-3},10^{-4}$, for fixed $\lambda/\lambda_{0}\simeq 1.33$ and $\lambda/\lambda_{0}\simeq 1.72$, and $T_{\rm c}=10\,\mathrm{MeV}$, $r_{T\rm,c}=0.4$. The blue contour marks the boundary of the physical region.
@@
Table 4: Spin and other physical properties of typical astrophysical objects.
We can see from Fig. 13 that a wide range of $s_{\rm rms,*}/(v_{\rm eff,*}/A_{\rm s})$ values can be obtained, even for fixed $A$ and $\lambda$, and for fixed $T_{\rm c}$ and $r_{T\rm,c}$. Equation (LABEL:MainFinalsRMSEstimate) explains this behavior because the percolation temperature and $\beta_{*}/H_{*}$ vary widely within the physical region. Furthermore, we see the effect of changing $T_{\rm c}$ and $r_{T\rm,c}$ in the contours of $s_{\rm rms,*}/(v_{\rm eff,*}/A_{\rm s})$ in Fig. 14. Panel (d) shows the scan for the default choice $T_{\rm c}=10\,\mathrm{MeV},r_{T\rm,c}=0.4$. Panels (a) and (b) show the effect of changing $T_{\rm c}$, which leads to increasing $s_{\rm rms,*}/(v_{\rm eff,*}/A_{\rm s})$ for larger $T_{\rm c}$ due to the increase in $g_{\rho,{\rm SM}}$, consistent with Eq. (LABEL:MainFinalsRMSEstimate). On the other hand, comparing panels (c) and (d) illustrates the effect of changing $r_{T\rm,c}$, where a smaller $r_{T\rm,c}$ increases the overall $s_{\rm rms,*}/(v_{\rm eff,*}/A_{\rm s})$.
@@
Figure 14: Contours of constant $\log_{10}\left[s_{\rm rms,*}/(v_{\rm eff,*}/A_{\rm s})\right]$ in steps of 0.5 for $A=10^{-3}$ and $\lambda/\lambda_{0}\simeq 1.33$, changing $T_{\rm c}$ (top) and $r_{T\rm,c}$ (bottom). The blue contour marks the boundary of the physical region.
@@ 5.3 Benchmarks
To obtain a general idea of the prediction for the RMS spin value of FV bubbles, we consider a collection of FOPT scenarios that will serve as our benchmark points. We require that these FOPT scenarios are, at least, physically viable. Then we calculate the transfer functions for these benchmarks, in order to eventually obtain $v_{\rm eff,*}$. One criterion we impose on these benchmarks is $T_{*}/T_{\rm c}\leq 0.95$, which guarantees that the effective potential will have a sufficiently high barrier between the false vacuum and the true vacuum. We also require that the cutoff scale $k_{\rm cut}$ be less than the scale $k_{\rm th}$ at which the RMS density contrast is $\simeq 0.4$; otherwise, we relax this condition so that $k_{\rm cut}/k_{\rm th}$ is at most $2$.
@@
The above requirements constrain the FOPT scenarios we feature to have $A=10^{-3},10^{-4}$; then for each $A$, we select $\lambda/\lambda_{0}(A)\simeq 1.33,1.72$. For each pair of $(A,\lambda)$, we scan the physical region in the ($L_{\rm c}/T_{\rm c}^{4},x_{C}$) plane and obtain physical quantities such as $\alpha_{*}$, $T_{*}$, and $s_{\rm rms,*}/(v_{\rm eff,*}/A_{\rm s})$. We choose a random collection of FOPT scenarios and solve the perturbation equations to construct a table of transfer functions versus $k/\mathcal{H}_{\rm c}$. This allows us to calculate $\mathcal{C}_{*}$ and the effective equatorial velocity $v_{\rm eff,*}/A_{\rm s}$, and finally $s_{\rm rms,*}$. The final benchmark points are presented in Table 1, with the rows organized in three groups. The first group (from the top) has parameters that specify the effective potential, critical temperature, and initial temperature ratio. The second group has quantities that can be obtained solely by tracking the evolution of the background cosmology during the phase transition. The third group has quantities that are relevant in inferring the features of the transfer functions, as well as quantities that can be obtained after solving the perturbation equations.
[…]
We now discuss the RMS spins for the benchmark points. We use Eq. (LABEL:MainFinalsRMSEstimate) as our guide to understand the dependence of $s_{\rm rms,*}/(v_{\rm eff,*}/A_{\rm s})$ — which we will call the RMS spin prefactor, since it excludes $v_{\rm eff,*}$ — on $T_{\rm c}$, $r_{T,\rm*}$, and $(\beta/H)_{*}/v_{\rm w,*}$. Firstly, the dependence on $T_{\rm c}$ enters in $g_{\rho,\rm SM}$, so we expect the benchmark with the lowest $T_{\rm c}$ (i.e., BP-2) to give the smallest RMS spin prefactor. Since all of the benchmarks have $v_{\rm w,*}$ close to unity, the RMS spin prefactor effectively scales with $(\beta/H)_{*}$. The RMS spin prefactor for BP-3 is close to that of BP-2 which has a lower $T_{\rm c}$, because the inverse FOPT time scale $(\beta/H)_{*}$ for BP-3 is lower than that for BP-2. Similarly, while BP-1 and BP-4 have the same $T_{\rm c}$, the inverse FOPT time scale for BP-4 is larger by about a factor of 5, so the RMS spin prefactor of BP-4 is larger by about a factor of 25. Comparing BP-3 and BP-5, which have the same $T_{\rm c}$ and roughly the same $(\beta/H)_{*}$, BP-5 has a $r_{T,\rm c}\approx r_{T\rm,*}\approx 0.1$. Since the RMS spin prefactor scales as $r_{T\rm,c}^{-4}$, it will be larger by $4^{4}=256$ for BP-5. Ultimately, the trend in $s_{\rm rms,*}$ is determined both by the RMS spin prefactor and $v_{\rm eff,*}$. The dependence of the latter on the FOPT scenario can only be determined once the transfer functions at percolation are known. With our analysis, we can only conclude that $v_{\rm eff,*}/A_{\rm s}$ can range from $\mathcal{O}(10^{-5})$ to $\mathcal{O}(10^{-3})$. For our benchmark points, the RMS spin ranges from $\mathcal{O}(10^{-5})$ to $\mathcal{O}(10)$.